### v1

# TESGNN: Temporal Equivariant Scene Graph Neural Networks for Efficient and Robust Multi-View 3D Scene Understanding

Abstract
Scene graphs have proven to be highly effective for various scene understanding tasks due to their compact and explicit representation of relational information. However, current methods often overlook the critical importance of preserving symmetry when generating scene graphs from 3D point clouds, which can lead to reduced accuracy and robustness, particularly when dealing with noisy, multi-view data. This work, to the best of our knowledge, presents the first implementation of an Equivariant Scene Graph Neural Network (ESGNN) to generate semantic scene graphs from 3D point clouds, specifically for enhanced scene understanding. Furthermore, a significant limitation of prior methods is the absence of temporal modeling to capture time-dependent relationships among dynamically evolving entities within a scene. To address this gap, we introduce a novel temporal layer that leverages the symmetry-preserving properties of ESGNN to fuse scene graphs across multiple sequences into a unified global representation by an approximate graph-matching algorithm. Our combined architecture, termed the Temporal Equivariant Scene Graph Neural Network (TESGNN), not only surpasses existing state-of-the-art methods in scene estimation accuracy but also achieves faster convergence. Importantly, TESGNN is computationally efficient and straightforward to implement using existing frameworks, making it well-suited for real-time applications in robotics and computer vision. This approach paves the way for more robust and scalable solutions to complex multi-view scene understanding challenges. Our source code is publicly available at: https://github.com/HySonLab/TESGraph.
1 Introduction
Figure 1: A visualization of multi-view scene graph from multiple 3D point cloud sequences. Our proposed TESGNN first generates local scene graphs for each sequence using Equivariant GNN. Then, the local scene graphs are merged by passing through a temporal layer to form a global scene graph representing the entire scene.
Holistic scene understanding plays a critical role in the advancement of robotics and computer vision systems [23, 29, 25]. The primary methodologies in this domain include 2D mapping, 3D reconstruction, and scene graph representation. Although 2D mapping remains a straightforward and widely adopted approach [37, 35], it inherently lacks a comprehensive spatial understanding of the environment. In contrast, 3D reconstruction techniques [20] provide richer spatial information but are resource intensive and particularly susceptible to noise interference [28]. Recent innovations, such as LMNav [46] and VLMap [14], leverage Large Language Models (LLMs) to generate semantic maps from 3D data, thereby enhancing scene understanding. However, these approaches often require substantial computational resources. As an alternative, scene graphs, particularly those utilizing Graph Neural Networks (GNNs), offer a more efficient and semantically rich representation compared to traditional 3D reconstruction techniques [6]. Initial methodologies mainly focused on generating scene graphs from sequences of 2D images. In more recent work [52, 53], researchers have explored scene graph generation by incorporating 3D features, such as depth data and point clouds. The latest advancements [54] propose hybrid approaches that integrate both 2D and 3D data to further refine scene graph representations. However, based on our analysis, existing approaches reveal two major limitations.
First, there is a noticeable lack of robust strategies for handling noisy, multi-view 3D data, which often results in inconsistencies within the generated scene graphs. These inconsistencies stem from the sensitivity of current methods to varying camera angles, ultimately failing to preserve symmetry. To address this, the symmetry-preserving property in Graph Neural Networks (GNNs) can be enhanced through: (1) employing invariant features for node and edge embeddings, and (2) leveraging Equivariant GNNs [26, 31]. Notably, prior works [53, 54] have yet to explore these symmetry-preserving techniques.
Second, existing methods lack effective mechanisms for constructing a coherent global scene graph from multiple point cloud sequences. For instance, [53] utilizes a heuristic averaging approach to merge several locally generated scene graphs into a single global representation. While conceptually straightforward, this method does not incorporate any optimization procedures, resulting in a global scene graph that is merely an unrefined aggregation of local graphs, lacking semantic coherence and structural consistency.
To address these limitations, we propose our novel architecture Temporal Equivariant Scene Graph Neural Network (TESGNN) that contains two main components:
  1. 1.
    Equivariant Scene Graph Neural Network (ESGNN): Starting from a point cloud input, ESGNN generates node and edge embeddings to construct a scene graph. To achieve this, we design an Equivariant GNN architecture that integrates the Equivariant Graph Convolution Layer [45] with a Feature-wise Attention layer [53]. This design ensures both rotational and translational invariance, thereby improving the quality of scene representation while reducing the need for deeper layers and extensive computational resources.
  2. 2.
    Temporal Model for Merging Scene Graphs: Leveraging the local scene graphs produced by the ESGNN, this Temporal Model fuses them by approximately solving a graph matching problem via comparing the similarity of output node embeddings. The resulting unified global scene graph can subsequently be utilized for tasks such as multi-agent coordination and efficient navigation in dynamic environments.
In summary, our paper presents two key contributions:
  • To the best of our knowledge, we are the first to apply Equivariant Graph Neural Networks for semantic scene graph generation from 3D point clouds. We empirically demonstrate that incorporating symmetry-preserving techniques significantly enhances the quality of scene graph representations. Our method achieves superior accuracy and requires fewer training iterations compared to existing approaches.
  • We introduce a novel embedding-based temporal approach to merge multiple scene graphs using embedding similarity and graph matching. This is, to our knowledge, the first of its kind, and we believe it will inspire further research into temporal architectures to refine global scene graphs to achieve improved holistic multi-view scene understanding.
2 Related Work
Graph Representation Learning.
Graph representation learning has emerged as a fundamental approach for encoding structural information into low-dimensional embeddings, enabling the application of machine learning techniques to graph-structured data. The primary objective of this field is to capture the topology of graphs along with node and edge attributes [36, 11, 39, 16], facilitating tasks such as node classification, link prediction, and graph classification. Works in this domain have introduced various methods, ranging from earlier spectral techniques to recent advancements in GNNs. Graph Transformers [56, 22, 4, 32] and Graph Attention Networks [50] extend the principles of traditional transformers and attention mechanisms to graph structures, enabling the capture of long-range interactions between nodes and enhancing the model’s ability to learn complex relationships in graph data. Graph representation learning is crucial for scene graph understanding as it efficiently captures the spatial relationships among entities, providing a compact and structured representation of the environment without the need for dense point cloud data.
Equivariant Neural Networks.
While Graph Neural Networks (GNNs) are inherently permutation-equivariant [17, 27], they do not inherently preserve rotational symmetry, which is crucial for accurately capturing 3D transformations such as translations, rotations, and reflections. Achieving rotational equivariance is particularly vital when working with 3D data formats like point clouds [49], 3D meshes [2], molecular structures [1, 40], and for Neural Relational Inference in systems with interacting entities [24]. Recent advancements have focused on enforcing SE(3) or E(3) symmetries in neural networks [47, 1, 10, 9, 26]. Notably, Satorras et al. [45] introduced the E(n)-equivariant GNN (EGNN), a powerful architecture that remains equivariant under translations, rotations, reflections, and permutations for sets of input points. Building on this, Liao and Smidt [31] developed the Equiformer, which enhances Transformer GNNs by replacing standard operations with equivariant counterparts, such as tensor products. In our work, we adopt the Equivariant Graph Convolution Layer from EGNN due to its demonstrated empirical effectiveness, integrating it into our framework for improved scene graph representation.
Scence Graph Understanding.
Research on scene graphs has significantly advanced tasks in vision, natural language processing, and interdisciplinary domains such as mobile robot navigation. Originally introduced as a structure to represent object instances and their relationships within a scene [21], scene graphs effectively capture rich semantics in various data modalities, including 2D/3D images [21, 3], videos [38], and point clouds [52]. Wald et al. [52] pioneered 3DSSG, the first method for generating 3D scene graphs from point clouds, accompanied by the annotated 3RScan dataset. Building on this, Wu et al. [53] introduced Scene Graph Fusion, which employs feature-wise attention to improve scene graph representations, followed by an incremental approach [54] that integrates both RGB image sequences and sparse 3D point clouds. Scene Graph Fusion remains the state-of-the-art method for scene graph generation from 3D point clouds. Our approach builds upon this by integrating Feature-wise Attention to optimize the Message-Passing process, thereby enhancing the quality of the generated scene graphs.
Temporal Graph Learning.
Temporal graph learning has gained significant attention in recent years, driven by the need to model dynamically evolving relationships over time with various applications in traffic prediction [55, 30, 34] and pandemic modeling [18, 48, 33]. Traditional Graph Neural Networks (GNNs) often assume static graph structures, which limits their applicability in scenarios where graph nodes and edges change continuously, such as in video analysis and multi-view scene understanding. To address these limitations, Temporal GNNs including [43, 44] incorporate temporal dependencies and time-series learning to capture time-evolving patterns and enhance predictive performance on dynamic graphs. Our work builds upon these concepts by introducing a temporal layer that preserves symmetry, leveraging the equivariant properties of our Equivariant Scene Graph Neural Network (ESGNN). Unlike prior approaches that treat temporal information as sequential snapshots, we propose a graph matching mechanism to merge scene graphs from multiple time steps into a unified global representation, which preserves temporal consistency while reducing computational overhead.
Large-scale Point Cloud Reconstruction.
Reconstructing large-scale point clouds is inherently challenging due to their irregular structure, often comprising multiple sub-point cloud sequences [15]. To tackle this issue, we introduce a temporal model that merges graphs generated from these sequences by utilizing embeddings from our proposed models. This approach not only improves integration across sequences but also extends to applications such as multi-agent SLAM and dynamic environment adaptation [19, 57], offering an efficient, lightweight solution based on multi-view scene graph representations.
3 Overall Framework
Fig. 2 illustrates the proposed framework’s capability to iteratively estimate a global 3D semantic scene graph using sequence of point clouds. The framework can be divided into two main parts, which are the feature extraction (a-c) mentioned in section 3.2 and scene graph generation (d-h) mentioned in 3.3.
Figure 2: Overview of our proposed Temporal Equivariant Scene Graph framework. Our approach takes sequences of point clouds a) as input to generate a geometric segmentations b). Subsequently, the properties of each segment and a neighbor graph between segments are constructed. The properties d) and neighbor graph e) of the segments that have been updated in the current frame c) are used as the inputs to compute node and edge features f) and to predict a 3D scene graph g). Then it goes through the temporal layer to fuse graphs from different sequences to a global one h).
3.1 Problem Formulation
Our system processes point cloud representations $D_{i}$ for each scene $i$ in a sequence $(D_{1},D_{2},\dots,D_{n})$ to generate corresponding scene graphs $(\mathcal{G}_{s,1},\mathcal{G}_{s,2},\dots,\mathcal{G}_{s,n})$ and a global graph $\mathcal{G}_{s,\text{global}}$ that aggregates scene information. Assuming effective feature extraction and geometric segmentation, our focus is on scene graph generation rather than feature extraction. Each point cloud $D_{i}$ undergoes pre-processing for input to a graph neural network (GNN). A geometric segmentation model partitions $D_{i}$ into segments, forming nodes of the scene graph, with segment attributes extracted via a point encoder. Detailed feature extraction methods are in Section 3.2.
We model symmetry preservation using the Euclidean group $E(3)$, which encompasses 3D rotations and translations. By employing equivariant layers, our model remains robust to variations in object orientation and position, enhancing generalization and performance while accelerating convergence in scene graph generation.
[…]
$\begin{cases}P(\mathcal{G}|D_{i})=P(\mathcal{G}|D_{j})_{i\neq j}\\ D_{j}=R_{i\rightarrow j}D_{i}+T_{i\rightarrow j}\end{cases}$ (1)
where $R_{i\rightarrow j}$ is the rotation matrix and $T_{i\rightarrow j}$ is the transition matrix.
Let us say we have a global point cloud, that is,
$D_{\text{global}}=\Phi_{\text{matching}}\left({D_{1},D_{2},...,D_{n}}\right),$
where $\Phi_{\text{matching}}$ is a model / algorithm to align subsequences together, we can define the global scene graph distribution as $P(\mathcal{G}_{s,\text{global}}|D_{\text{global}})$. However, $\Phi_{\text{matching}}$ is hard to define, especially in the case that the origin coordinate of each sequence is not aligned, which consumes time and resources for sampling or large estimation models. Instead, as the symmetry-perserving property is maintained, the embedding between each graph is similar, it is much easier to perform the matching between the graph embedding vectors, so that:
$P(\mathcal{G}_{s,\text{global}}|D_{\text{global}})$ $=P(\mathcal{G}_{s}|\Phi_{\text{matching}}(\left({D_{1},D_{2},...,D_{n}}\right))$
$=\Phi_{\text{graph}}\left(\mathcal{G}_{s,1},\mathcal{G}_{s,2},...,\mathcal{G}_{s,n}\right)$
where $\Phi_{graph}$ is the model matching the embedding vectors together, which will be described in the later section.
3.2 Feature Extraction
In this phase, the framework extracts the feature for scene graph generation, following the two main steps:
  1. 1.
    Point Cloud Reconstruction: The proposed framework takes the point cloud data, which can be reconstructed from various techniques such as ORB-SLAM3 or HybVIO [5] as the input. However, for the objective validation purpose of scene graph generation, we use the indoor point cloud dataset 3RScan [51] for ground truth data $D_{i}$.
  2. 2.
    Geometric Segmentation and Point Cloud Extraction with Segment Nodes: Given a point cloud $D_{i}$, this geometric segmentation will provide a segment set $\mathbf{S}=\left\{\mathbf{s}_{1},\ldots,\mathbf{s}_{n}\right\}$. Each segment $s_{i}$ consists of a set of 3D points $\mathbf{P}_{i}$ where each point is defined as a 3D coordinate $p_{i}\in\mathbb{R}^{3}$ and RGB color. Then, the point cloud concerning each entity is fed to the point encoder named PointNet [7] $f_{p}(\mathbf{P_{i}})$ to encode the segments $s_{i}$ into latent node and edge features, which are then passed to the model detailed in Section 4.
3.3 Scene Graph Generation
In this phase, the framework processes the input from feature extraction (Section 3.2) to generate the scene graph:
  • Properties and Neighbor Graph Extraction: From the point cloud, we extract features including the centroid $\overline{\mathbf{p}}_{i}\in\mathbb{R}^{3}$, standard deviation $\boldsymbol{\sigma}_{i}\in\mathbb{R}^{3}$, bounding box size $\mathbf{b}_{i}\in\mathbb{R}^{3}$, maximum length $l_{i}\in\mathbb{R}$, and volume $\nu_{i}\in\mathbb{R}$. We create edges between nodes only if their bounding boxes are within 0.5 meters of each other, following [53].
  • Scene Graph Representation and Temporal Graph Matching: The processed input from above is then fed to the ESGNN (Section 4) to generate the node and edge embeddings. These embeddings are used to generate the sub-graph for each sequence and merge multiple scene graphs with our Temporal Model (Section 5). For ESGNN, the node classes and edge predicates are predicted using two Multi-Layer Perceptron (MLP) classifiers. The network is trained end-to-end with a joint cross-entropy loss for both objects $\mathcal{L}_{\text{obj}}$ and predicates $\mathcal{L}_{\text{pred}}$, as described in [52]. On the other hand, our Temporal Model is trained to minimize the representation distance between identical nodes with contrastive loss from [12].
Figure 3: Overview of the ESGNN (Equivariant Scene Graph Neural Network) Architecture. The model comprises two main layers: (1) Feature-wise Attention Graph Convolution Layer (FAN-GCL) and (2) Equivariant Graph Convolution Layer (EGCL). FAN-GCL handles large inputs with multi-head attention to update node and edge features, while EGCL ensures symmetry preservation by incorporating bounding box coordinates into the message-passing mechanism. ESGNN leverages these layers to maintain rotation and permutation equivariance, and enhance the quality of scene graph generation from 3D point clouds while reducing computational costs.
4 Equivariant Scene Graph Neural Network (ESGNN)
To build a network architecture that effectively generate scene graph from point cloud, we propose a combination of Equivariant Graph Convolution Layers [45] and the Graph Convolution Layers with Feature-wise Attention [53].
4.1 Graph Initialization
Node features
The node feature includes the invariant features $\mathbf{h_{i}}$ and the 3-vector coordinate $\mathbf{x_{i}}\in\mathbb{R}^{3}$. The invariant features $\mathbf{h}_{i}$ consists of the latent feature of the point cloud after going through the PointNet $f_{p}(\mathbf{P_{i}})$, standard deviation $\boldsymbol{\sigma}_{i}$, log of the bounding box size $\ln\left(\mathbf{b}_{i}\right)$ where $\mathbf{b}_{i}=\left(b_{x},b_{y},b_{z}\right)\in\mathbb{R}^{3}$, log of the bounding box volume $\ln(v_{i})$ where $v_{i}=b_{x}b_{y}b_{z}$, and log of the maximum length of bounding box $\ln(l_{i})$. The coordinate of the bounding box $x_{i}$ is defined by the coordinates of the two furthest corners of the bound box. The formula can be written as follows:
$\mathbf{v}_{i}$ $=\left(\mathbf{h}_{i},\mathbf{x}_{i}\right),$
$\mathbf{h}_{i}$ $=\left[f_{p}\left(\mathbf{P}_{i}\right),\boldsymbol{\sigma}_{i},\ln\left(\mathbf{b}_{i}\right),\ln\left(\nu_{i}\right),\ln\left(l_{i}\right)\right],$
$\mathbf{x}_{i}$ $=[\mathbf{x}_{i}^{\text{bottom right}},\mathbf{x}_{i}^{\text{top left}}].$
Edge features
The visual characteristics of the edges are determined by the properties of the connected segments. For an edge between a source node $i$ and a target node $j$ where $j\neq i$, the edge visual feature $\mathbf{e}_{ij}$ is computed as follows:
$\begin{gathered}\mathbf{r}_{ij}=\left[\overline{\mathbf{p}}_{i}-\overline{\mathbf{p}}_{j},\boldsymbol{\sigma}_{i}-\boldsymbol{\sigma}_{j},\mathbf{b}_{i}-\mathbf{b}_{j},\ln\left(\frac{l_{i}}{l_{j}}\right),\ln\left(\frac{\nu_{i}}{\nu_{j}}\right)\right],\\ \mathbf{e}_{ij}=g_{s}\left(\mathbf{r}_{ij}\right),\end{gathered}$
where $g_{s}(\cdot)$ is a Multi-Layer Perceptron (MLP) that projects the paired segment properties into a latent space.
4.2 Graph Neural Network Architecture and Weights Updating
Fig. 3 describes our GNN network with two core components: ① Feature-wise Attention Graph Convolution Layer (FAN-GCL); and ② Equivariant Graph Convolution Layer (EGCL). The FAN-GCL, proposed by [53], is used to handle the large input queries $Q$ of dimensions $d_{q}$ and targets $T$ of dimensions $d_{\tau}$ by utilizing multi-head attention. On the other hand, EGCL, proposed by [45], is used to maintain symmetry-preserving equivariance, allowing us to incorporate the bounding box coordinates $x_{i}$ as node features and update them through the message-passing mechanism.
ESGNN has 2 main layers, each consists of a FAN-GCL followed by an EGCL, forming a total of 4 message-passing layers. For each main layer, the formulas for updating node and edge features are defined as:
  • Update FAN-GCL:
    $\mathbf{v}_{i}^{\ell+1}$ $=g_{v}\left(\left[\mathbf{v}_{i}^{\ell},\max_{j\in\mathcal{N}(i)}\left(\operatorname{FAN}\left(\mathbf{v}_{i}^{\ell},\mathbf{e}_{ij}^{\ell},\mathbf{v}_{j}^{\ell}\right)\right)\right]\right),$
    $\mathbf{e}_{ij}^{\ell+1}$ $=g_{e}\left(\left[\mathbf{v}_{i}^{\ell},\mathbf{e}_{ij}^{\ell},\mathbf{v}_{j}^{\ell}\right]\right),$
  • Update EGCL:
    $h_{i}^{(l+1)}$ $=h_{i}^{(l)}+\text{g}_{\text{v}}\left(\text{concat}\left(h_{i}^{(l)},\sum_{j\in\mathcal{N}(i)}e_{ij}^{(l)}\right)\right),$
    $e_{ij}^{(l+1)}$ $=\text{g}_{\text{e}}\left(\text{concat}\left(h_{i}^{(l)},h_{j}^{(l)},\|\mathbf{x}_{i}^{(l)}-\mathbf{x}_{j}^{(l)}\|^{2},e_{ij}^{(l)}\right)\right),$
    $\mathbf{x}_{i}^{(l+1)}$ $=\mathbf{x}_{i}^{(l)}+\sum_{j\in\mathcal{N}(i)}(\mathbf{x}_{i}^{(l)}-\mathbf{x}_{j}^{(l)})\cdot\phi_{\text{coord}}(e_{ij}^{(l)}).$
5 Temporal Model for Merging Scene Graphs
Figure 4: Temporal Matching Pipeline. The process begins with node and edge embeddings derived from different sequences (a). For each sequence, edge embeddings are concatenated with the target node embeddings to create Predicate-Object embeddings (b), which then pass through a linear layer followed by sum pooling (c). For each node, the embeddings from all associated edges are concatenated and processed through another linear layer and self-attention mechanism to generate the final embedding representation for each segment (d). These final embeddings are utilized for top K graph matching (e).
By leveraging the symmetry-preserving properties of our equivariant network, we hypothesize that its node and edge embeddings can remain inherently distinguishable, especially for isomorphic scene graphs subjected to rotations and translations. This unique feature opens the gate for an interesting application: Given sequences of point clouds capturing overlapping regions of the same scene, our model can reliably match the corresponding parts, regardless of viewpoint changes. As a result, it facilitates the merging of multiple scene graphs into a unified global representation, ensuring spatial consistency and completeness over time.
Figure 4 illustrates our ESGNN pipeline for graph matching and merging. Since sequences may differ in the number of nodes and edges, we define a Node-Edge representation for each node and its connected edges. Our goal is to produce consistent representations for identical node-edge embeddings appearing across sequences.
[…]
Contrastive Loss Our training goal is to optimize the embedding model such that distances between similar node-edge embeddings are minimized, while those between dissimilar embeddings are maximized. We achieve this using Contrastive Loss [12], a well-established method for training embeddings [41]. Unlike traditional loss functions that aggregate over samples, Contrastive Loss operates on pairs:
$\mathcal{L}(W,Y,X_{1},X_{2})=\frac{1-Y}{2}(D_{W})^{2}+\frac{Y}{2}\max(0,m-D_{W})^{2},$
where $(Y,X_{1},X_{2})$ are pairs of node-edge embeddings, $D_{W}$ is a distance function, and $m>0$ is a margin. We use Siamese Cosine Similarity [42] for $D_{W}$ and train with Online Contrastive Loss, focusing on hard positive and hard negative pairs. These are widely-used methods in Sentence Transformer [41], which inspire our approach for the Temporal Model.
6 Experiment
(a) Edge recall on evaluation set 3DSSG-l20
(b) Edge recall on evaluation set 3DSSG-l20
(c) Edge recall on evaluation set 3DSSG-l20
(d) Node recall on evaluation set 3DSSG-l20
(e) Node recall on evaluation set 3DSSG-l20
(f) Node recall on evaluation set 3DSSG-l20
Figure 5: (Comparison of recall result over training steps, column-wise interpretation. (a), (d) illustrates the result of our model versus SGFN. (b), (e) shows the ablation study. (c), (f) shows result while applying our model with the image encoder.
We evaluate TESGNN on the 3DSSG dataset [52] and compare the results with state-of-the-art works. Section 6.2 provides our model performance for scene graph generation from 3D point clouds data, in comparison to 3DSSG [52] and Scene Graph Fusion (SGFN) [53]. Our method is evaluated on full scenes given geometric segments mentioned in section 3.2. Section 6.4 discusses an extended version of our model incorporating both 2D image sequences and point clouds, built upon JointSSG [54]. Section 6.5 reports our Temporal Model for graph matching.
6.1 Dataset and Metrics
Dataset: We use the 3DSSG [52] - a dataset for scene graph generation built upon 3RScan[51], adapting the setting from [53]  [^1] https://github.com/ShunChengWu/3DSSG. The scene graph ground truths in 3DSSG are divided into 2 versions: l20, which includes 20 objects and 8 predicates, and l160, which includes 160 objects and 26 predicates. We mainly use the test set of the l20 version for our experiment as the l160 is currently having some issues within the dataset.. The original 3RScan contains 1482 3D reconstructions/snapshots of 478 naturally changing indoor environments. After processed with ScanNet [8] for geometric segmentation and grouth truth scene graph generation, the final dataset consists of 1061 sequences from 382 scenes for train, 157 sequences from 47 scenes for validation, and 117 sequences from 102 scenes for test.
[…]
Additional metrics: We assess our Temporal Model by following the Information Retrieval evaluation metrics, including R@k, mAP@k, and MRR@k. mAP@k measures the quality of rankings by emphasizing $k$ relevant nodes at higher ranks. MRR@k evaluates the effectiveness of retrieval by averaging the reciprocal ranks of the $k$ similar nodes. We also report the resource used for training and inference.
6.2 Scene Graph Generation Evaluation
Superior prediction performance: Table 1 compares the results between ESGNN (ours) with existing models 3DSSG [52] and SGFN [53] on the 3DSSG-l20 dataset with geometric segmentation setting. Our method obtains high results in both relationship, object, and predicate classification. Especially, ESGNN outperforms the existing methods in relationship prediction and obtains significantly higher R@1 in predicate compared to SGFN. ESGNN also works well with unseen data, with competitive result compared to SGFN, as shown in Table 2.
Method Relationship Object Predicate Recall
R@1 R@3 R@1 R@3 R@1 R@2 Obj. Rel.
3DSSG 32.65 50.56 55.74 83.89 $\mathbf{95.22}$ 98.29 55.74 $\mathbf{95.22}$
SGFN 37.82 48.74 62.82 $\mathbf{88.08}$ 81.41 98.22 63.98 94.24
Ours $\mathbf{43.54}$ $\mathbf{53.64}$ $\mathbf{63.94}$ 86.65 94.62 $\mathbf{98.30}$ $\mathbf{65.45}$ 94.62
Table 1: Scene graph predictions for relationship triplet, object, and predicate, measured on 3DSSG-l20. The Recall column reports the recall scores on objects (Obj.) and relationships (Rel.)
Method New Relationship New Object New Predicate
R@1 R@3 R@1 R@3 R@1 R@2
3DSSG 39.74 49.79 55.89 84.42 $\mathbf{70.87}$ 83.29
SGFN $\mathbf{47.01}$ 55.30 64.50 $\mathbf{88.92}$ 68.71 $\mathbf{83.76}$
Ours 46.85 $\mathbf{56.95}$ $\mathbf{65.47}$ 87.52 66.90 82.88
Table 2: Scene graph predictions for new unseen relationship triplet, object, and predicate, measured on 3DSSG-l20 with geometric segmentation.
Faster training convergence: Figure 5(a) and 5(d) report the recalls for nodes (objects) and edges (relationships) during training between ESGNN and SGFN on both train and validation sets. The recall slope of ESGNN in the first 10 epochs (5000 steps) is significantly higher than that of SGFN. This indicates that ESGNN has faster convergence and higher initial recall.
(a) ESGNN edge
(b) SGFN edge
(c) ESGNN node
(d) SGFN node
Figure 6: Edge and node confusion matrices after the first 1000 steps (2 epochs). ESGNN is better at handling imbalanced classes, as the predictions spread across multiple classes rather than “blindly” predicting towards dominant classes like SGFN.
More robust against imbalanced data: We additionally provide the training confusion matrices after the first 1000 steps in Figure 6. It is notably observable that ESGNN 6(a), 6(c) shows clearer diagonal patterns compared to SGFN 6(b), 6(d), which was biased towards some dominant classes. These patterns demonstrate that ESGNN is more data-efficient than SGFN, as it does not need to generalize over rotations and translations of the data, while still harnessing the flexibility of GNNs in larger datasets. However, due to its high bias [45], our EGCL-based message passing performs well with limited data but struggles to learn the dataset’s subtleties as the training size increases. In later training steps, we observe the overfitting. Yet, our model consistently outperforms the pioneering works overall.
6.3 Ablation Study
Table 3 evaluates of ESGNN with different architectures and settings that we experimented. ⓪ is the SGFN, run as the baseline model for comparison. ① is the ESGNN with 1 FAN-GCL layer and 1 EGCL layer. This is our best performer and is used for experiments in Section 6.2. ② is similar to ①. The only difference is that we concatenate coordinate embedding to the output edge embedding after message passing. Meanwhile, ③ and ④ are similar to ① and ② respectively, but with 2 FAN-GCL layers and 2 EGCL layers.
Method Relationship Object Predicate
R@1 R@3 R@1 R@3 R@1 R@2
⓪ SGFN 37.82 48.74 62.82 $\mathbf{88.08}$ 81.41 98.22
ESGNN_1 $\mathbf{42.30}$ $\mathbf{53.30}$ $\mathbf{63.21}$ 86.70 94.34 $\mathbf{98.30}$
② ESGNN_1X 34.96 42.59 57.55 86.18 92.68 98.08
③ ESGNN_2 35.63 44.63 57.55 84.41 93.93 97.94
④ ESGNN_2X 37.94 50.58 59.97 85.23 $\mathbf{94.53}$ 98.01
Table 3: Evaluation of different ESGNN architectures on scene graph generation task on 3DSSG-l20 dataset. ② is our best performer and is used for the evaluation in Section 6.2.
Models ③ and ④ perform well on the train set but poorly on the validation and test sets, potentially suffering overfitting as they contain more layers. Models ② and ④ have higher edge recalls in several initial epochs, but experience a decline in the later epochs, as shown in Fig 5(b), 5(e).
6.4 ESGNN with Image Encoder
Our model also poses a potential in application on point-image encoders model together such as JointSSG [54]. We implement our GNN architecture similar to JointSSG and name it Joint-ESGNN. Figures 5(c) and 5(f) show the performance of our model with image encoders compared to JointSSG and SGFN.
6.5 Temporal Graph Matching Evaluation
[…]
Metric@k k=1 k=2 k=3 k=5
R@k 0.7015 0.8406 0.9009 0.9525
mAP@k 0.7015 0.7893 0.8194 0.8409
MRR@k 0.7015 0.7711 0.7911 0.8032
Table 4: Evaluation for our Temporal Model with 10 training epochs. Notice that k=1 is the accuracy of the node-matching.
Prediction performance: Our Temporal Model is trained on 321,344 pairs of node with 16,161 positive pairs generated from the 3DSSG dataset. Table 4 demonstrates its strong accuracy in node-matching, highlighting its ability to effectively identify similar nodes.
Potential Customization and Improvement: As graph-matching for point cloud scene graphs is a novel approach, no existing methods are available for direct comparison. Given our model’s embedding-based nature, there are various opportunities for improvement. For instance, incorporating multiple loss functions, such as the Multiple Negatives Ranking Loss [13] alongside Contrastive Loss, could significantly enhance training performance. Various data sampling strategies could also be explored to enhance the model’s robustness.
6.6 Run time
We measure the runtime of our system during the training phase, which demands the most resources. Our machine is equipped with two Intel Xeon Gold 6242 CPUs, each running at 3.9 GHz with a total of 64 threads, and four Nvidia GeForce RTX 3090 GPUs. However, we utilized only 8 data workers during training, resulting in a total CPU utilization of 0.3%. Only one GPU was used, consuming a maximum of 1.6 GB of memory per experiment. The average training time per step, excluding data loading, is approximately 0.25 seconds. Inference time is significantly shorter, and our model’s size is only 37 MB.
7 Conclusion
In this paper, we introduced the TESGNN, a novel method that overcomes key limitations in 3D scene understanding. By leveraging the symmetry-preserving property of the Equivariant GNN, our architecture ensures robust and efficient generation of semantic scene graphs from 3D point clouds. Our proposed Temporal Model integrates multiple scene graphs, providing a global representations for real-time dynamic environments. Experimental results show that TESGNN outperforms state-of-the-art methods in accuracy, convergence speed, and computational efficiency.
Limitations & Future Work
Future work will focus on optimizing TESGNN for complex real-world scenarios by integrating additional sensor modalities like LiDAR and RGB-D data to improve scene graph generation. We also aim to extend the model’s temporal capabilities for continuous data streams, enhancing its suitability for autonomous navigation and multi-agent systems.