### v2

# TESGNN: Temporal Equivariant Scene Graph Neural Networks for Efficient and Robust Multi-View 3D Scene Understanding

Abstract
Scene graphs have proven to be highly effective for various scene understanding tasks due to their compact and explicit representation of relational information. However, current methods often overlook the critical importance of preserving symmetry when generating scene graphs from 3D point clouds, which can lead to reduced accuracy and robustness, particularly when dealing with noisy, multi-view data. Furthermore, a major limitation of prior approaches is the lack of temporal modeling to capture time-dependent relationships among dynamically evolving entities in a scene. To address these challenges, we propose Temporal Equivariant Scene Graph Neural Network (TESGNN), consisting of two key components: (1) an Equivariant Scene Graph Neural Network (ESGNN), which extracts information from 3D point clouds to generate scene graph while preserving crucial symmetry properties, and (2) a Temporal Graph Matching Network, which fuses scene graphs generated by ESGNN across multiple time sequences into a unified global representation using an approximate graph-matching algorithm. Our combined architecture TESGNN outperforms current state-of-the-art methods in scene graph generation, achieving higher accuracy and faster training convergence. Moreover, we show that leveraging the symmetry-preserving property produces a more stable and accurate global scene representation compared to existing approaches. Last but not least, it is computationally efficient and easily implementable using existing frameworks, making it well-suited for real-time applications in robotics and computer vision. This approach paves the way for more robust and scalable solutions to complex multi-view scene understanding challenges. Our source code is publicly available at: https://github.com/HySonLab/TESGraph.
I Introduction
Figure 1: A visualization of multi-view scene graph from multiple 3D point cloud sequences. Our proposed TESGNN first generates local scene graphs for each sequence using Equivariant GNN. Then, the local scene graphs are merged by passing through a temporal layer to form a global scene graph representing the entire scene.
Holistic scene understanding plays a critical role in the advancement of robotics and computer vision systems [1, 2, 3]. The primary methodologies in this domain include 2D mapping, 3D reconstruction, and scene graph representation. Although 2D mapping remains a straightforward and widely adopted approach [4], it inherently lacks a comprehensive spatial understanding of the environment. In contrast, 3D reconstruction techniques [5] provide richer spatial information but are resource intensive and particularly susceptible to noise interference [6]. Recent innovations, such as LMNav [7] and VLMap [8], leverage Large Language Models (LLMs) to generate semantic maps from 3D data, thereby enhancing scene understanding. However, these approaches often require substantial computational resources. As an alternative, scene graphs, particularly those utilizing Graph Neural Networks (GNNs), offer a more efficient and semantically rich representation compared to traditional 3D reconstruction techniques [9]. Initial methodologies mainly focused on generating scene graphs from sequences of 2D images. In more recent work [10, 11], researchers have explored scene graph generation by incorporating 3D features, such as depth data and point clouds. The latest advancements [12] propose hybrid approaches that integrate both 2D and 3D data to further refine scene graph representations. However, based on our analysis, existing approaches reveal two major limitations.
Current methods struggle with noisy, multi-view 3D data, leading to inconsistent scene graphs due to sensitivity to camera angles and poor symmetry preservation. Enhancing Graph Neural Networks (GNNs) with invariant node and edge embeddings and using Equivariant GNNs [13, 14] can address this, though prior works [11, 12] haven’t explored these approaches. Additionally, existing methods lack robust strategies for building coherent global scene graphs from multiple point cloud sequences. For example, [11] uses heuristic averaging to merge local graphs but skips optimization, resulting in global graphs that lack semantic coherence and structural consistency.
We introduce the Temporal Equivariant Scene Graph Neural Network (TESGNN), a novel architecture that combines two key innovations:
  • Equivariant Scene Graph Neural Network (ESGNN) for Scene Graph Extraction: extracts scene graphs from 3D point clouds using an Equivariant GNN with Equivariant Graph Convolution and Feature-wise Attention layers, ensuring rotational and translational invariance. This design delivers high-quality scene representations with superior accuracy, fewer training iterations, and reduced computational demands compared to existing methods—marking the first application of symmetry-preserving Equivariant GNNs for this task.
  • Temporal Scene Graph Matching for Global Scene Representation: fuses the local scene graphs into a unified global representation by solving a graph matching problem based on node embedding similarity, with ESGNN as the backbone scene graph extractor. This embedding-based temporal approach achieves robust multi-view scene understanding, outperforming other extractors and enabling applications like multi-agent coordination and navigation in dynamic environments.
II Related Work
Graph Representation and Equivariance
Graph representation learning encodes structural and relational data into low-dimensional embeddings using methods like GNNs, Graph Transformers [15, 16, 17], and Graph Attention Networks [18], enabling tasks such as scene graph understanding without dense 3D data. For 3D applications, standard GNNs lack rotational symmetry; recent SE(3)/E(3)-equivariant architectures like EGNN [19] and Equiformer [14] address this by preserving geometric symmetries during transformations, critical for point clouds [20], molecular structures, and spatial reasoning.
Scence Graph Understanding
Research on scene graphs has significantly advanced tasks in vision, natural language processing, and interdisciplinary domains such as mobile robot navigation. Originally introduced as a structure to represent object instances and their relationships within a scene [21], scene graphs effectively capture rich semantics in various data modalities, including 2D/3D images [21, 22], videos [23], and point clouds [10]. [10] pioneered 3DSSG, the first method for generating 3D scene graphs from point clouds, accompanied by the annotated 3RScan dataset. Building on this, [11] introduced Scene Graph Fusion, which employs feature-wise attention to improve scene graph representations, followed by an incremental approach [12] that integrates both RGB image sequences and sparse 3D point clouds. Scene Graph Fusion remains the state-of-the-art method for scene graph generation from 3D point clouds. Our approach builds upon this by integrating Feature-wise Attention to optimize the Message-Passing process, thereby enhancing the quality of the generated scene graphs.
Temporal Graph Learning
Temporal graph learning has gained significant attention for modeling dynamic relationships over time, with applications in traffic prediction [24, 25, 26]. Traditional GNNs assume static graph structures, limiting their applicability in scenarios like video analysis and multi-view scene understanding. To address this, Temporal GNNs [27, 28] incorporate time-series learning to capture evolving patterns. In scene graph matching, methods such as SGAligner [29] and SG-PGM [30] align nodes across graphs but rely on ground truth graphs, making them impractical for real-world tasks where predicted graphs contain noise, ambiguous edges, and node permutations. To overcome these challenges, we introduce a symmetry-preserving temporal layer in ESGNN, leveraging equivariant properties to merge scene graphs across time steps. Unlike prior approaches that treat temporal information as sequential snapshots, our method integrates graph matching to construct a unified representation that maintains temporal consistency while reducing computational overhead.
Figure 2: Overview of our proposed TESGNN. Our approach takes sequences of point clouds a) as input to generate a geometric segmentations b). Subsequently, the properties of each segment and a neighbor graph between segments are constructed. The properties d) and neighbor graph e) of the segments that have been updated in the current frame c) are used as the inputs to compute node and edge features f) and to predict a 3D scene graph g). Then it goes through the temporal layer to fuse graphs from different sequences to a global one h).
Large-scale Point Cloud Reconstruction
Reconstructing large-scale point clouds is inherently challenging due to their irregular structure, often comprising multiple sub-point cloud sequences [31]. To tackle this issue, we introduce a temporal model that merges graphs generated from these sequences by utilizing embeddings from our proposed models. This approach not only improves integration across sequences but also extends to applications such as multi-agent SLAM and dynamic environment adaptation [32, 33], offering an efficient, lightweight solution based on multi-view scene graph representations.
III Overall Framework
Fig. 2 illustrates the proposed framework’s capability to iteratively estimate a global 3D semantic scene graph from sequence of point clouds. The framework consists of three key phases: Feature Extraction (a-c), discussed in Section III-B; Scene Graph Extraction (d-f); and Temporal Graph Matching (g-h), detailed in Section III-C. Our main works, scene graph extraction and temporal graph matching, are further elaborated in Sections IV and V.
III-A Problem Formulation
Our system processes point cloud representations $D_{i}$ for each scene $i$ in a sequence $(D_{1},D_{2},\dots,D_{n})$ to generate corresponding scene graphs $(\mathcal{G}_{s,1},\mathcal{G}_{s,2},\dots,\mathcal{G}_{s,n})$ and a global graph $\mathcal{G}_{s,\text{global}}$ that aggregates scene information. Assuming effective feature extraction and geometric segmentation, our focus is on scene graph generation rather than feature extraction. Each point cloud $D_{i}$ undergoes pre-processing for input to a graph neural network (GNN). A geometric segmentation model partitions $D_{i}$ into segments, forming nodes of the scene graph, with segment attributes extracted via a point encoder. Detailed feature extraction methods are in Section III-B.
We model symmetry preservation using the Euclidean group $E(3)$, which encompasses 3D rotations and translations. By employing equivariant layers, our model remains robust to variations in object orientation and position, enhancing generalization and performance while accelerating convergence in scene graph generation.
[…]
$\begin{cases}P(\mathcal{G}|D_{i})=P(\mathcal{G}|D_{j})_{i\neq j}\\ D_{j}=R_{i\rightarrow j}D_{i}+T_{i\rightarrow j}\end{cases}$ (1)
where $R_{i\rightarrow j}$ is the rotation matrix and $T_{i\rightarrow j}$ is the transition matrix.
Let us say we have a global point cloud, that is,
$D_{\text{global}}=\Phi_{\text{matching}}\left({D_{1},D_{2},...,D_{n}}\right),$
where $\Phi_{\text{matching}}$ is a model / algorithm to align subsequences together, we can define the global scene graph distribution as $P(\mathcal{G}_{s,\text{global}}|D_{\text{global}})$. However, $\Phi_{\text{matching}}$ is hard to define, especially in the case that the origin coordinate of each sequence is not aligned, which consumes time and resources for sampling or large estimation models. Instead, as the symmetry-perserving property is maintained, the embedding between each graph is similar, it is much easier to perform the matching between the graph embedding vectors, so that:
$P(\mathcal{G}_{s,\text{global}}|D_{\text{global}})$ $=P(\mathcal{G}_{s}|\Phi_{\text{matching}}(\left({D_{1},D_{2},...,D_{n}}\right))$
$=\Phi_{\text{graph}}\left(\mathcal{G}_{s,1},\mathcal{G}_{s,2},...,\mathcal{G}_{s,n}\right)$
where $\Phi_{graph}$ is the model matching the embedding vectors together, which will be described in the later section.
III-B Feature Extraction
In this phase, the framework extracts the feature for scene graph generation, following the two main steps:
  1. 1.
    Point Cloud Reconstruction: The proposed framework takes the point cloud data, which can be reconstructed from various techniques such as ORB-SLAM3 or HybVIO [34] as the input. However, for the objective validation purpose of scene graph generation, we use the indoor point cloud dataset 3RScan [35] for ground truth data $D_{i}$.
  2. 2.
    Geometric Segmentation and Point Cloud Extraction with Segment Nodes: Given a point cloud $D_{i}$, this geometric segmentation will provide a segment set $\mathbf{S}=\left\{\mathbf{s}_{1},\ldots,\mathbf{s}_{n}\right\}$. Each segment $s_{i}$ consists of a set of 3D points $\mathbf{P}_{i}$ where each point is defined as a 3D coordinate $p_{i}\in\mathbb{R}^{3}$ and RGB color. Then, the point cloud concerning each entity is fed to the point encoder named PointNet [36] $f_{p}(\mathbf{P_{i}})$ to encode the segments $s_{i}$ into latent node and edge features, which are then passed to the model detailed in Section IV.
III-C Scene Graph Extraction and Temporal Graph Matching
In this phase, the framework processes the input from feature extraction (Section III-B) to generate the scene graph:
  • Properties and Neighbor Graph Extraction: From the point cloud, we extract features including the centroid $\overline{\mathbf{p}}_{i}\in\mathbb{R}^{3}$, standard deviation $\boldsymbol{\sigma}_{i}\in\mathbb{R}^{3}$, bounding box size $\mathbf{b}_{i}\in\mathbb{R}^{3}$, maximum length $l_{i}\in\mathbb{R}$, and volume $\nu_{i}\in\mathbb{R}$. We create edges between nodes only if their bounding boxes are within 0.5 meters of each other, following [11].
  • Scene Graph Extraction and Temporal Graph Matching: The processed input from above is then fed to the ESGNN (Section IV) to generate the node and edge embeddings. These embeddings are used to generate the sub-graph for each sequence and merge multiple scene graphs with our Temporal Model (Section V). For ESGNN, the node classes and edge predicates are predicted using two Multi-Layer Perceptron (MLP) classifiers. ESGNN is trained end-to-end with a joint cross-entropy loss for both objects $\mathcal{L}_{\text{obj}}$ and predicates $\mathcal{L}_{\text{pred}}$, as described in [10]. Meanwhile, our Temporal Graph Matching is trained to minimize the representation distance between identical nodes with contrastive loss [37].
Figure 3: ESGNN Scene Graph Extractor pipeline. The model comprises two main layers: (1) Feature-wise Attention Graph Convolution Layer (FAN-GCL) and (2) Equivariant Graph Convolution Layer (EGCL). FAN-GCL handles large inputs with multi-head attention to update node and edge features, while EGCL ensures symmetry preservation by incorporating bounding box coordinates into the message-passing mechanism. ESGNN leverages these layers to maintain rotation and permutation equivariance, thus enhance the quality of scene graph generation.
IV Equivariant Scene Graph Neural Network (ESGNN)
To build a network architecture that effectively generate scene graph from point cloud, we propose a combination of Equivariant Graph Convolution Layers [19] and the Graph Convolution Layers with Feature-wise Attention [11].
IV-A Graph Initialization
Node features
The node feature includes the invariant features $\mathbf{h_{i}}$ and the 3-vector coordinate $\mathbf{x_{i}}\in\mathbb{R}^{3}$. The invariant features $\mathbf{h}_{i}$ consists of the latent feature of the point cloud after going through the PointNet $f_{p}(\mathbf{P_{i}})$, standard deviation $\boldsymbol{\sigma}_{i}$, log of the bounding box size $\ln\left(\mathbf{b}_{i}\right)$ where $\mathbf{b}_{i}=\left(b_{x},b_{y},b_{z}\right)\in\mathbb{R}^{3}$, log of the bounding box volume $\ln(v_{i})$ where $v_{i}=b_{x}b_{y}b_{z}$, and log of the maximum length of bounding box $\ln(l_{i})$. The coordinate of the bounding box $x_{i}$ is defined by the coordinates of the two furthest corners of the bound box. The formula can be written as follows:
$\mathbf{v}_{i}$ $=\left(\mathbf{h}_{i},\mathbf{x}_{i}\right),$
$\mathbf{h}_{i}$ $=\left[f_{p}\left(\mathbf{P}_{i}\right),\boldsymbol{\sigma}_{i},\ln\left(\mathbf{b}_{i}\right),\ln\left(\nu_{i}\right),\ln\left(l_{i}\right)\right],$
$\mathbf{x}_{i}$ $=[\mathbf{x}_{i}^{\text{bottom right}},\mathbf{x}_{i}^{\text{top left}}].$
Edge features
The visual characteristics of the edges are determined by the properties of the connected segments. For an edge between a source node $i$ and a target node $j$ where $j\neq i$, the edge visual feature $\mathbf{e}_{ij}$ is computed as follows:
$\begin{gathered}\mathbf{r}_{ij}=\left[\overline{\mathbf{p}}_{i}-\overline{\mathbf{p}}_{j},\boldsymbol{\sigma}_{i}-\boldsymbol{\sigma}_{j},\mathbf{b}_{i}-\mathbf{b}_{j},\ln\left(\frac{l_{i}}{l_{j}}\right),\ln\left(\frac{\nu_{i}}{\nu_{j}}\right)\right],\\ \mathbf{e}_{ij}=g_{s}\left(\mathbf{r}_{ij}\right),\end{gathered}$
where $g_{s}(\cdot)$ is a Multi-Layer Perceptron (MLP) that projects the paired segment properties into a latent space.
IV-B Graph Neural Network Architecture and Weights Updating
Fig. 3 describes our GNN network with two core components: ① Feature-wise Attention Graph Convolution Layer (FAN-GCL); and ② Equivariant Graph Convolution Layer (EGCL). The FAN-GCL, proposed by [11], is used to handle the large input queries $Q$ of dimensions $d_{q}$ and targets $T$ of dimensions $d_{\tau}$ by utilizing multi-head attention. On the other hand, EGCL, proposed by [19], is used to maintain symmetry-preserving equivariance, allowing us to incorporate the bounding box coordinates $x_{i}$ as node features and update them through the message-passing mechanism.
ESGNN has 2 main layers, each consists of a FAN-GCL followed by an EGCL, forming a total of 4 message-passing layers. For each main layer, the formulas for updating node and edge features are defined as:
  • Update FAN-GCL:
    $\mathbf{v}_{i}^{\ell+1}$ $=g_{v}\left(\left[\mathbf{v}_{i}^{\ell},\max_{j\in\mathcal{N}(i)}\left(\operatorname{FAN}\left(\mathbf{v}_{i}^{\ell},\mathbf{e}_{ij}^{\ell},\mathbf{v}_{j}^{\ell}\right)\right)\right]\right),$
    $\mathbf{e}_{ij}^{\ell+1}$ $=g_{e}\left(\left[\mathbf{v}_{i}^{\ell},\mathbf{e}_{ij}^{\ell},\mathbf{v}_{j}^{\ell}\right]\right),$
  • Update EGCL:
    $h_{i}^{(l+1)}$ $=h_{i}^{(l)}+\text{g}_{\text{v}}\left(\text{concat}\left(h_{i}^{(l)},\sum_{j\in\mathcal{N}(i)}e_{ij}^{(l)}\right)\right),$
    $e_{ij}^{(l+1)}$ $=\text{g}_{\text{e}}\left(\text{concat}\left(h_{i}^{(l)},h_{j}^{(l)},\|\mathbf{x}_{i}^{(l)}-\mathbf{x}_{j}^{(l)}\|^{2},e_{ij}^{(l)}\right)\right),$
    $\mathbf{x}_{i}^{(l+1)}$ $=\mathbf{x}_{i}^{(l)}+\sum_{j\in\mathcal{N}(i)}(\mathbf{x}_{i}^{(l)}-\mathbf{x}_{j}^{(l)})\cdot\phi_{\text{coord}}(e_{ij}^{(l)}).$
V Temporal Graph Matching Network
Figure 4: Temporal Graph Matching pipeline. First, node and edge embeddings derived from scene graphs of different sequences (a). For each sequence, edge embeddings are concatenated with the target node embeddings to create Predicate-Object embeddings (b), which then pass through a linear layer followed by sum pooling (c). For each node, the embeddings from all associated edges are concatenated and processed through another linear layer and self-attention mechanism to generate the final representation of each segment (d). These final embeddings are utilized for top-K retrieval graph matching (e).
By leveraging the symmetry-preserving properties of our equivariant network, we hypothesize that its node and edge embeddings can remain inherently distinguishable, especially for isomorphic scene graphs subjected to rotations and translations. This unique feature opens the gate for an interesting application: Given sequences of point clouds capturing overlapping regions of the same scene, our model can reliably match the corresponding parts, regardless of viewpoint changes. As a result, it facilitates the merging of multiple scene graphs into a unified global representation, ensuring spatial consistency and completeness over time.
Figure 4 illustrates our graph matching. Since sequences may differ in the number of nodes and edges, we define a Node-Edge representation for each node and its connected edges. Our goal is to produce consistent representations for identical node-edge embeddings appearing across sequences.
[…]
Contrastive Loss Our training goal is to optimize the embedding model such that distances between similar node-edge embeddings are minimized, while those between dissimilar embeddings are maximized. We achieve this using Contrastive Loss [37], a well-established method for training embeddings [38]. Unlike traditional loss functions that aggregate over samples, Contrastive Loss operates on pairs:
$\mathcal{L}(W,Y,X_{1},X_{2})=\frac{1-Y}{2}(D_{W})^{2}+\frac{Y}{2}\max(0,m-D_{W})^{2},$
where $(Y,X_{1},X_{2})$ are pairs of node-edge embeddings, $D_{W}$ is a distance function, and $m>0$ is a margin. We use Siamese Cosine Similarity [38] for $D_{W}$ and train with Online Contrastive Loss, focusing on hard positive and hard negative pairs. These are widely-used methods in Sentence Transformer [38], which inspire our approach.
(a) Edge recall on evaluation set 3DSSG-l20
(b) Edge recall on evaluation set 3DSSG-l20
(c) Edge recall on evaluation set 3DSSG-l20
(d) Node recall on evaluation set 3DSSG-l20
(e) Node recall on evaluation set 3DSSG-l20
(f) Node recall on evaluation set 3DSSG-l20
Figure 5: Comparison of recall result over training steps, column-wise interpretation. (a), (d) illustrates the result of our model versus SGFN. (b), (e) shows the ablation study. (c), (f) shows result while applying our model with the image encoder.
VI Experiment
We evaluate TESGNN on the 3DSSG dataset [10] and compare the results with state-of-the-art works. Section VI-C provides result for scene graph generation from 3D point clouds, in comparison to 3DSSG [10] and Scene Graph Fusion (SGFN) [11]. Our method is evaluated on full scenes given geometric segments mentioned in Section III-B. Section VI-E reports our Temporal Graph Matching.
VI-A Dataset
Scene Graph Extraction: We use the 3DSSG [10]  [^1] https://github.com/ShunChengWu/3DSSG. The scene graph ground truth in 3DSSG contains 2 versions: l20, which includes 20 objects and 8 predicates, and l160, which includes 160 objects and 26 predicates. We mainly use the test set of the l20 version for our experiment as the l160 is currently having some issues within the dataset, as confirmed by the authors. - a popular dataset for 3D scene graph evaluation built upon 3RScan[35], adapting the setting from SGFN [11]. The original 3RScan contains 1482 3D reconstructions/snapshots of 478 naturally changing indoor environments. After processed with ScanNet [39] for geometric segmentation and ground truth scene graph generation, the final dataset consists of 1061 sequences from 382 scenes for training, 157 sequences from 47 scenes for validation, and 117 sequences from 102 scenes for test.
Temporal Graph Matching: While reproducing prior works including SGAligner [29] and SG-PGM [30], we identified limitations in their dataset. Since their work does not contain scene graph extraction, they synthesize scene graph node and edge features using one-hot encoding of the ground truth labels, which does not accurately reflect scene graph embeddings. This set-up impractically ignores real-world noises affecting local scene graphs. To address these shortcomings, we leverage our above set-up for scene graph extraction to evaluate scene graph matching, using the same sequences and scenes. In this context, a positive pair consists of the same object appearing in two or more different sequences. There are 16,156 positive pairs over 321,292 total pairs for train set, and 3,062 positive pairs over 63,890 total pairs for test set. In our set-up, node and edge features are derived from frozen scene graphs such as ESGNN (ours) or SGFN [11].
VI-B Metrics
[…]
VI-C Scene Graph Extraction Evaluation
Better prediction outcome: Table I compares the results between ESGNN (ours) with existing models 3DSSG [10] and SGFN [11] on the 3DSSG-l20 dataset with geometric segmentation setting. Our method obtains high results in both relationship, object, and predicate classification. Especially, ESGNN outperforms the existing methods in relationship prediction and obtains significantly higher R@1 in predicate compared to SGFN. For unseen data, ESGNN is competitive with SGFN, as shown in Table II.
Method Relationship Object Predicate Recall
R@1 R@3 R@1 R@3 R@1 R@2 Obj. Rel.
3DSSG 32.65 50.56 55.74 83.89 $\mathbf{95.22}$ 98.29 55.74 $\mathbf{95.22}$
SGFN 37.82 48.74 62.82 $\mathbf{88.08}$ 81.41 98.22 63.98 94.24
Ours $\mathbf{43.54}$ $\mathbf{53.64}$ $\mathbf{63.94}$ 86.65 94.62 $\mathbf{98.30}$ $\mathbf{65.45}$ 94.62
TABLE I: Scene graph predictions for relationship triplet, object, and predicate, measured on 3DSSG-l20. The Recall column reports the recall scores on objects (Obj.) and relationships (Rel.)
Method New Relationship New Object New Predicate
R@1 R@3 R@1 R@3 R@1 R@2
3DSSG 39.74 49.79 55.89 84.42 $\mathbf{70.87}$ 83.29
SGFN $\mathbf{47.01}$ 55.30 64.50 $\mathbf{88.92}$ 68.71 $\mathbf{83.76}$
Ours 46.85 $\mathbf{56.95}$ $\mathbf{65.47}$ 87.52 66.90 82.88
TABLE II: Scene graph predictions for new unseen relationship triplet, object, and predicate, measured on 3DSSG-l20 with geometric segmentation.
Faster training convergence: Figure 5(a) and 5(d) report the recalls for nodes (objects) and edges (relationships) during training of ESGNN and SGFN on both train and validation sets. The recall slope of ESGNN in the first 10 epochs (5000 steps) is significantly higher than that of SGFN. This shows that ESGNN has faster convergence and higher initial recall.
(a) ESGNN edge
(b) SGFN edge
(c) ESGNN node
(d) SGFN node
Figure 6: Confusion matrices after the first 1000 steps (2 epochs). ESGNN is better at handling imbalanced classes, as the predictions spread across multiple classes rather than “blindly” predicting towards dominant classes like SGFN.
More robust against imbalanced data: We additionally provide the training confusion matrices after the first 1000 steps in Figure 6. It is notably observable that ESGNN 6(a), 6(c) shows clearer diagonal patterns compared to SGFN 6(b), 6(d), which was biased towards some dominant classes. These patterns demonstrate that ESGNN is more data-efficient than SGFN, as it does not need to generalize over rotations and translations of the data, while still harnessing the flexibility of GNNs in larger datasets. However, due to its high bias [19], our EGCL-based message passing performs well with limited data but struggles to learn the dataset’s subtleties as the training size increases. In later training steps, we observe the overfitting. Yet, our model consistently outperforms the pioneering works overall.
Scene Graph Extractor with Image Encoder: ESGNN are also potential to apply with point-image encoder instead of 3D point clouds only. Figures 5(c) and 5(f) show the performance of our implementation with image encoders, called Joint-ESGNN, compared to existing methods including JointSSG [12] and SGFN.
VI-D Ablation Study
Table III evaluates of ESGNN with different architectures and settings that we experimented. ⓪ is the SGFN, run as the baseline model for comparison. ① is the ESGNN with 1 FAN-GCL layer and 1 EGCL layer. This is our best performer and is used for experiments in Section VI-C. ② is similar to ①. The only difference is that we concatenate coordinate embedding to the output edge embedding after message passing. ③ and ④ are similar to ① and ② respectively, but with 2 FAN-GCL layers and 2 EGCL layers.
Method Relationship Object Predicate
R@1 R@3 R@1 R@3 R@1 R@2
⓪ SGFN 37.82 48.74 62.82 $\mathbf{88.08}$ 81.41 98.22
ESGNN_1 $\mathbf{42.30}$ $\mathbf{53.30}$ $\mathbf{63.21}$ 86.70 94.34 $\mathbf{98.30}$
② ESGNN_1X 34.96 42.59 57.55 86.18 92.68 98.08
③ ESGNN_2 35.63 44.63 57.55 84.41 93.93 97.94
④ ESGNN_2X 37.94 50.58 59.97 85.23 $\mathbf{94.53}$ 98.01
TABLE III: Evaluation of different ESGNN architectures on scene graph generation task on 3DSSG-l20 dataset. ② is our best performer and is used for evaluation in Section VI-C.
Models ③ and ④ perform well on the train set but poorly on the validation and test sets, potentially suffering overfitting as they contain more layers. Models ② and ④ have higher edge recalls in several initial epochs, but decline in the later epochs, as shown in Fig 5(b), 5(e).
VI-E Temporal Graph Matching Evaluation
Experiment Setup: We assess our work from two key perspectives. First, we evaluate our Graph Matching Network using different backbone scene graph extractors, including the state-of-the-art SGFN [11] and our proposed ESGNN. Second, we compare the effectiveness of the SGFN and ESGNN backbones on the state-of-the-art SG-PGM [30]. To ensure the same set-up, we reimplement and evaluate SG-PGM with our proposed dataset mentioned in Section VI-A.
Method Backbone R@k MRR@k
@1 @2 @3 @5 @2 @3
SG-PGM SGFN 10.61 19.16 28.56 44.71 14.88 18.02
ESGNN $\mathbf{30.14}$ $\mathbf{46.94}$ $\mathbf{57.63}$ $\mathbf{73.86}$ $\mathbf{38.54}$ $\mathbf{42.10}$
SG-PGM SGFN 18.54 33.28 $\mathbf{45.73}$ $\mathbf{62.82}$ 25.91 30.06
(point) ESGNN $\mathbf{20.91}$ $\mathbf{33.55}$ 44.41 62.01 $\mathbf{27.23}$ $\mathbf{30.85}$
Ours SGFN 54.48 72.95 82.11 90.45 63.71 66.77
ESGNN $\mathbf{70.15}$ $\mathbf{84.06}$ $\mathbf{90.09}$ $\mathbf{95.25}$ $\mathbf{78.93}$ $\mathbf{81.94}$
TABLE IV: Evaluation for Graph Matching Network with 30 training epochs. Notice that k=1 is the accuracy of the node-matching and R@1 = MRR@1
Better outcome with ESGNN backbone scene graph extractor: Table IV demonstrates that utilizing our proposed ESGNN as the backbone scene graph extractor consistently outperforms SGFN across all metrics, on both graph matching networks. This underscores the advantage of ESGNN’s symmetry-preserving property, which enhances the representation of nodes and edges in 3D point clouds, making them more resilient to transformations and significantly improving graph matching performance.
Better outcome with our Graph Matching Network: Although SG-PGM [30] achieved competitive results for node matching in their setting, it performs poorly in our scenario, where node and edge features are extracted from a scene graph extractor model rather than one-hot encoded from the ground truth. In our case, both nodes and edges have more labels, leading to several ambiguous predictions, including overlaps, incorrect assignments, and permutations. This explains why SG-PGM fails in this setting. In contrast, our model leverages the equivariant properties of nodes and edges while applying sum pooling, ensuring permutation invariance and significantly outperforming SG-PGM in this scenario, indicating the robustness of the model.
VII Conclusion
In this paper, we introduced the TESGNN, a novel method that overcomes key limitations in 3D scene understanding. By leveraging the symmetry-preserving property of the Equivariant GNN, our architecture ensures robust and efficient generation of semantic scene graphs from 3D point clouds. Our proposed Temporal Graph Matching Model provides a global representation from local scene graphs for real-time dynamic environments. Experimental results show that TESGNN outperforms state-of-the-art methods in accuracy, convergence speed, and computational efficiency.
Limitations & Future Work Future work will focus on optimizing TESGNN for complex real-world scenarios by integrating additional sensor modalities like LiDAR and RGB-D data to improve scene graph generation. We also aim to extend the model’s temporal capabilities for continuous data streams, enhancing its suitability for autonomous navigation and multi-agent systems.